OpenStudy (anonymous):

Solve and graph the absolute value inequality: |2x + 1| ≤ 5. number line with closed dots on -3 and 2 with shading going in the opposite directions. number line with closed dots on -3 and 2 with shading in between. number line with open dots on -3 and 2 with shading in between. number line with closed dots on -2 and 2 with shading in between.

OpenStudy (texaschic101):

| 2x + 1 | <= 5 -5 <= 2x + 1 <= 5 --- subtract 1 -1 - 5 <= 2x <= 5 - 1 -6 <= 2x <= 4 --- divide by 2 -3 <= x <= 2 so basically this is saying that x is greater or equal to -3, bu less or equal to 2 if there is an equal sign, the it has a closed circle < less then, means shaded to the left > greater then, means shaded to the right solution : closed circle on -3, closed circle on 2, shaded in between